Reserve Problems Chapter 11 Section 2 Problem 1 The department of health studied the number of patients who need liver transpType or paste question here
Answer #1

X Υ Xy χ2 y2 2005 1355 2716775.000 4020025.000 1836025.000 2006 1794 3218436.000 3598764.000 5926671.000 4024036.000 4028049.The slope B1, B1 = (n*xx y.- (2x)*(y))/(n*Zx2 - (2x)) B1 = (13*160728589.000 - (26143.000)*(79863.000))/(13*52573755.000 - (2a) ση2 = Σy2-βΟΣ»-β1Σχy 1-2 584432131--1365049.308):79863-681.846160728589 13-2 = 768293.375 b) β1 =681.846 c) Χ=2020 Υ= 681.